OpenStudy (anonymous):

If 6 cards are drawn at random from a standard deck of 52 cards, what is the probability that exactly 2 of the cards are spades?

OpenStudy (anonymous):

(A).038, (B).200, (C).315, (D).465, (E).747

OpenStudy (anonymous):

@paki @thomaster

OpenStudy (paki):

i think 2/52 = 0.038...

OpenStudy (cwrw238):

THERE MIGHT BE AN EASier way to do this but here goes suppose the first 2 cards are spades and the last 4 are not spades the probability of this is (1/4)*(12/51)*(39/50)*(38/49)*(37/48)*(36/47) = 0.021 now there are 6C2 ways of combining 2 spades in 6 cards = 15 so the rquired probability = 15 *021 = 0.315

OpenStudy (cwrw238):

yes thats it

OpenStudy (anonymous):

Thank you @cwrw238 i knew there was something else to it.

OpenStudy (anonymous):

Thank you as well @paki

OpenStudy (cwrw238):

yw